# Euclid Quick Data Release (Q1)

### v1

@@
Euclid Collaboration: L. Bazzanini    G. Angora    P. Bergamini    M. Meneghetti    P. Rosati    A. Acebron    C. Grillo    M. Lombardi    R. Ratta    M. Fogliardi    G. Di Rosa    D. Abriola    M. D’Addona    G. Granata    L. Leuzzi    A. Mercurio    S. Schuldt    E. Vanzella    C. Tortora    B. Altieri    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    A. Balestra    S. Bardelli    P. Battaglia    A. Biviano    E. Branchini    M. Brescia    S. Camera    G. Cañas-Herrera    V. Capobianco    C. Carbone    J. Carretero    M. Castellano    G. Castignani    S. Cavuoti    A. Cimatti    C. Colodro-Conde    G. Congedo    L. Conversi    Y. Copin    A. Costille    F. Courbin    H. M. Courtois    M. Cropper    A. Da Silva    H. Degaudenzi    G. De Lucia    H. Dole    F. Dubath    C. A. J. Duncan    X. Dupac    S. Dusini    S. Escoffier    M. Fabricius    M. Farina    R. Farinelli    F. Faustini    S. Ferriol    F. Finelli    M. Frailis    E. Franceschi    M. Fumana    S. Galeotta    K. George    W. Gillard    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    J. Hoar    W. Holmes    I. M. Hook    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    A. Kiessling    M. Kilbinger    B. Kubik    M. Kunz    H. Kurki-Suonio    R. Laureijs    A. M. C. Le Brun    D. Le Mignant    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    G. Mainetti    D. Maino    E. Maiorano    O. Mansutti    O. Marggraf    M. Martinelli    N. Martinet    F. Marulli    R. J. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    C. Neissner    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    W. J. Percival    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    A. Renzi    J. Rhodes    G. Riccio    E. Romelli    M. Roncarelli    R. Saglia    Z. Sakr    A. G. Sánchez    D. Sapone    B. Sartoris    P. Schneider    T. Schrabback    A. Secroun    G. Seidel    S. Serrano    P. Simon    C. Sirignano    G. Sirri    L. Stanco    J. Steinwagner    P. Tallada-Crespí    A. N. Taylor    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    E. A. Valentijn    L. Valenziano    J. Valiviita    T. Vassallo    G. Verdoes Kleijn    A. Veropalumbo    Y. Wang    J. Weller    A. Zacchei    G. Zamorani    E. Zucca    M. Ballardini    M. Bolzonella    E. Bozzo    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    W. G. Hartley    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. Pezzotta    M. Pöntinen    I. Risso    V. Scottez    M. Sereno    M. Tenti    M. Viel    M. Wiesmann    Y. Akrami    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    E. Aubourg    D. Bertacca    M. Bethermin    A. Blanchard    L. Blot    H. Böhringer    M. Bonici    S. Borgani    M. L. Brown    S. Bruton    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    T. Castro    B. Clément    F. Cogato    S. Conseil    A. R. Cooray    O. Cucciati    S. Davini    F. De Paolis    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    P. Dimauro    P.-A. Duc    M. Y. Elkhashab    A. Enia    Y. Fang    A. Finoguenov    A. Fontana    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    R. Gavazzi    E. Gaztanaga    F. Giacomini    F. Gianotti    A. H. Gonzalez    G. Gozaliasl    M. Guidi    C. M. Gutierrez    S. Hemmati    H. Hildebrandt    J. Hjorth    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    J. Kim    C. C. Kirkpatrick    S. Kruk    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    T. I. Liaudat    S. J. Liu    A. Loureiro    J. Macias-Perez    M. Magliocchetti    F. Mannucci    R. Maoli    C. J. A. P. Martins    L. Maurin    C. J. R. McPartland    M. Miluzio    P. Monaco    C. Moretti    G. Morgante    C. Murray    K. Naidoo    A. Navarro-Alsina    S. Nesseris    D. Paoletti    F. Passalacqua    K. Paterson    A. Pisani    D. Potter    S. Quai    M. Radovich    P.-F. Rocci    S. Sacquegna    M. Sahlén    D. B. Sanders    E. Sarpa    A. Schneider    D. Sciotti    E. Sellentin    L. C. Smith    J. G. Sorce    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    D. Vergani    G. Verza    P. Vielzeuf    N. A. Walton
(November 4, 2025)
@@ 1 Introduction
The visual inspection of Euclid Q1 (Euclid Quick Release Q1 2025) cluster lens candidates carried out in Euclid Collaboration: Bergamini et al. (2025a) involved the coordinated effort of around $40$ expert astronomers to assess approximately $1300$ candidate clusters within large $4^{\prime}\times 4^{\prime}$ cutouts. This task required several weeks of distributed work, underscoring the resource-intensive nature of such visual inspection. While effective, this approach is not sustainable for future data releases, which will contain orders of magnitude more data; our goal with this work is precisely to address this challenge. This underscores the need for more efficient, scalable, and automated approaches to analyse the vast amounts of data produced by these surveys. In the last decade, several alternative methods have been developed for detecting SL events, particularly on galaxy-scale. These methods range from semi-automated algorithms designed to detect arc and ring-like structures (e.g. More et al. 2012; Gavazzi et al. 2014; Sonnenfeld et al. 2018) to crowd-sourced science initiatives (e.g. Marshall et al. 2016; Sonnenfeld et al. 2020). Within this framework, machine learning (ML) and deep learning (DL) techniques have emerged as the most promising tools for identifying SL events (e.g. Metcalf et al. 2019).
@@
The challenges of gravitational arc detection in galaxy clusters are well-suited for the Mask R-CNN framework; this advanced neural network (NN) indeed offers key advantages over standard CNNs. In particular, it can process full-size images of galaxy clusters thanks to its ability to handle inputs of variable size, without requiring resizing to a fixed input dimension, thereby preserving morphological features and spatial information. Moreover, it enables simultaneous detection, classification, and segmentation of multiple objects in a single image at different scales, allowing it to distinguish individual arc instances within crowded fields. These capabilities make Mask R-CNNs an effective choice for addressing the complexity of the cluster-scale lensing regime, positioning this work as an innovative application of advanced computer vision techniques to SL in wide-field surveys.
@@
This paper is organised as follows. In Sect. 2 we describe how the training set is produced via SL simulations. In Sect. 3, we introduce the Mask R-CNN framework, describing in detail the architecture of our implementation, and the training setup. In Sect. 4, we present the results of the trained network, we evaluate its performance on the test set, and we further test it on real Euclid Q1 images. Finally, in Sect. 5 we summarise, discuss the results, and conclude.
@@ 3.1 Implementation
Transfer learning is a technique that enables NNs to leverage knowledge acquired from one task to enhance performance on a related but distinct task (refer to Tan et al. 2018 for an overview). This methodology capitalises on pre-existing weight values derived from one data set, employing them as the foundation for network initialisation for training the network on another data set. The advantages of this technique are twofold: it accelerates the training process, and it mitigates network overfitting. In our work, we utilise as a starting point the Mask R-CNN weights pre-trained on the Microsoft Common Objects in Context (MS COCO; Lin et al. 2014) data set. The MS COCO data set comprises approximately 328 000328\,000 images, categorised into $91$ classes of everyday objects. After initialising the pre-trained NN parameters, we made trainable all its $\approx$ 44 million parameters, including also the backbone ones.
@@ 4.2 Performance on the test set
Figure 5: In panel (a) we show two sets of three PPRR curves for $\mathrm{IoU}_{\mathrm{thr}}=0.5,0.75$, @5050:55:$95$, each obtained by varying the object score threshold $p_{\mathrm{thr}}$. In panel (b) we show the PPRR curve for $\mathrm{IoU}_{\mathrm{thr}}=0.5$, where the dots are colour-coded by the score threshold $p_{\mathrm{thr}}$ used to evaluate them. In particular, we emphasise the PPRR pair associated to our choice of score threshold $p_{\mathrm{thr}}=0.996$. Note that the recall remains relatively high even at moderate precision, reflecting the network’s ability to recover most bright arcs while keeping the number of false positives manageable. Increasing the threshold shifts the operating point toward higher precision at the expense of recall. In panel (c) we show two precision and the recall curves as a function of the score threshold for $\mathrm{IoU}_{\mathrm{thr}}=0.5$.
@@ Acknowledgements.
We acknowledge financial support through grant 2020SKSTHZ. LB is indebted to the communities behind the multiple free, libre, and open-source software packages on which we all depend. MM was supported by INAF Grants The Big-Data era of cluster lensing” and ”Probing Dark Matter and Galaxy Formation in Galaxy Clusters through Strong Gravitational Lensing”, and ASI Grant n. 2024-10-HH.0 ”Attivit scientifiche per la missione Euclid fase E”. The research activities described in this paper have been co-funded by the European Union NextGeneration EU within PRIN 2022 project no. 20229YBSAN Globular clusters in cosmological simulations and in lensed fields: from their birth to the present epoch. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This work used the following software packages: Python (Van Rossum & Drake 2009), PyTorch (Paszke et al. 2019; Ansel et al. 2024), torchvision (TorchVision 2016), NumPy (van der Walt et al. 2011; Harris et al. 2020), SciPy (Virtanen et al. 2020) scikit-learn (Pedregosa et al. 2011) Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), matplotlib (Hunter 2007), lenstool (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009), ds9 (Joye & Mandel 2003), Aladin (Bonnarel et al. 2000), topcat (Taylor 2005), git, bash (GNU 2007).

### v2

@@
Euclid Collaboration: L. Bazzanini    G. Angora    P. Bergamini    M. Meneghetti    P. Rosati    A. Acebron    C. Grillo    M. Lombardi    R. Ratta    M. Fogliardi    G. Di Rosa    D. Abriola    M. D’Addona    G. Granata    L. Leuzzi    A. Mercurio    J. M. Palencia    S. Schuldt    E. Vanzella    C. Tortora    B. Altieri    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    A. Balestra    S. Bardelli    P. Battaglia    A. Biviano    E. Branchini    M. Brescia    S. Camera    G. Cañas-Herrera    V. Capobianco    C. Carbone    J. Carretero    M. Castellano    G. Castignani    S. Cavuoti    A. Cimatti    C. Colodro-Conde    G. Congedo    L. Conversi    Y. Copin    A. Costille    F. Courbin    H. M. Courtois    M. Cropper    A. Da Silva    H. Degaudenzi    G. De Lucia    H. Dole    F. Dubath    C. A. J. Duncan    X. Dupac    S. Dusini    S. Escoffier    M. Fabricius    M. Farina    R. Farinelli    F. Faustini    S. Ferriol    F. Finelli    M. Frailis    E. Franceschi    M. Fumana    S. Galeotta    K. George    W. Gillard    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    J. Hoar    W. Holmes    I. M. Hook    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    A. Kiessling    M. Kilbinger    B. Kubik    M. Kunz    H. Kurki-Suonio    R. Laureijs    A. M. C. Le Brun    D. Le Mignant    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    G. Mainetti    D. Maino    E. Maiorano    O. Mansutti    O. Marggraf    M. Martinelli    N. Martinet    F. Marulli    R. J. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier Deceased    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    C. Neissner    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    W. J. Percival    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    A. Renzi    J. Rhodes    G. Riccio    E. Romelli    M. Roncarelli    R. Saglia    Z. Sakr    A. G. Sánchez    D. Sapone    B. Sartoris    P. Schneider    T. Schrabback    A. Secroun    G. Seidel    S. Serrano    P. Simon    C. Sirignano    G. Sirri    L. Stanco    J. Steinwagner    P. Tallada-Crespí    A. N. Taylor    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    E. A. Valentijn    L. Valenziano    J. Valiviita    T. Vassallo    G. Verdoes Kleijn    A. Veropalumbo    Y. Wang    J. Weller    A. Zacchei    G. Zamorani    E. Zucca    M. Ballardini    M. Bolzonella    E. Bozzo    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    W. G. Hartley    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. Pezzotta    M. Pöntinen    I. Risso    V. Scottez    M. Sereno    M. Tenti    M. Viel    M. Wiesmann    Y. Akrami    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    E. Aubourg    D. Bertacca    M. Bethermin    A. Blanchard    L. Blot    H. Böhringer    M. Bonici    S. Borgani    M. L. Brown    S. Bruton    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    T. Castro    B. Clément    F. Cogato    S. Conseil    A. R. Cooray    O. Cucciati    S. Davini    F. De Paolis    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    P. Dimauro    P.-A. Duc    M. Y. Elkhashab    A. Enia    Y. Fang    A. Finoguenov    A. Fontana    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    R. Gavazzi    E. Gaztanaga    F. Giacomini    F. Gianotti    A. H. Gonzalez    G. Gozaliasl    M. Guidi    C. M. Gutierrez    S. Hemmati    H. Hildebrandt    J. Hjorth    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    J. Kim    C. C. Kirkpatrick    S. Kruk    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    T. I. Liaudat    S. J. Liu    A. Loureiro    J. Macias-Perez    M. Magliocchetti    F. Mannucci    R. Maoli    C. J. A. P. Martins    L. Maurin    C. J. R. McPartland    M. Miluzio    P. Monaco    C. Moretti    G. Morgante    C. Murray    K. Naidoo    A. Navarro-Alsina    S. Nesseris    D. Paoletti    F. Passalacqua    K. Paterson    A. Pisani    D. Potter    S. Quai    M. Radovich    P.-F. Rocci    S. Sacquegna    M. Sahlén    D. B. Sanders    E. Sarpa    A. Schneider    D. Sciotti    E. Sellentin    L. C. Smith    J. G. Sorce    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    D. Vergani    G. Verza    P. Vielzeuf    N. A. Walton    D. Scott
@@ 1 Introduction
The visual inspection of Euclid Q1 (Euclid Quick Release Q1 2025) cluster lens candidates carried out in Euclid Collaboration: Bergamini et al. (2025a) involved the coordinated effort of around $40$ expert astronomers to assess approximately $1300$ candidate clusters within large $4^{\prime}\times 4^{\prime}$ cutouts. This task required several weeks of distributed work, underscoring the resource-intensive nature of such visual inspection. While effective, this approach is not sustainable for future data releases, which will contain orders of magnitude more data; our goal with this work is precisely to address this challenge. This underscores the need for more efficient, scalable, and automated approaches to analyse the vast amounts of data produced by these surveys. In the last decade, several alternative methods have been developed for detecting SL events, particularly on galaxy-scale. These methods range from semi-automated algorithms designed to detect arc and ring-like structures (e.g. More et al. 2012; Gavazzi et al. 2014; Sonnenfeld et al. 2018) to crowd-sourced science initiatives (e.g. Marshall et al. 2016; Sonnenfeld et al. 2020). Several arc-finding algorithms proposed in the literature identify gravitational arcs by detecting elongated structures in the surface brightness distribution, for example by correlating local ellipticities across overlapping image cells (Seidel & Bartelmann 2007), or by segmenting objects using local intensity gradients and applying selection criteria based on elongation (e.g. length-to-width ratio) and morphological properties (Xu et al. 2016). Within this framework, machine learning (ML) and deep learning (DL) techniques have emerged as the most promising tools for identifying SL events (e.g. Metcalf et al. 2019).
@@
The challenges of gravitational arc detection in galaxy clusters are well-suited for the Mask R-CNN framework; this advanced neural network (NN) indeed offers key advantages over standard CNNs. In particular, it can process full-size images of galaxy clusters thanks to its ability to handle inputs of variable size, without requiring resizing to a fixed input dimension, thereby preserving morphological features and spatial information. We adopt Mask R-CNN not for pixel-level arc measurements, but because its instance-segmentation framework naturally enables simultaneous detection, classification, and segmentation of multiple objects in a single image at different scales, allowing it to distinguish individual arc instances within large, crowded fields. Our primary goal in this work is therefore lens finding on cluster scales, not obtaining high-precision masks for e.g. downstream automated lens modelling. These capabilities make Mask R-CNNs an effective choice for addressing the complexity of the cluster-scale lensing regime, positioning this work as an innovative application of advanced computer vision techniques to SL in wide-field surveys.
@@
This paper is organised as follows. In Sect. 2 we describe how the training set is produced via SL simulations. In Sect. 3, we introduce the Mask R-CNN framework, describing in detail the architecture of our implementation, and the training setup. In Sect. 4, we present the results of the trained network, we evaluate its performance on the test set, and we further test it on real Euclid Q1 images. In Sect. 5, we discuss the results. Finally, in Sect. 6 we summarise and conclude.
@@ 2.4 Building the knowledge base
Although the training set is based on only $10$ clusters, each of them provides a very rich lensing environment: they contain numerous real arcs and highly structured mass distributions, and by injecting simulated sources we can generate thousands of independent realisations that preserve realistic cluster morphologies.
@@ 3.1 Implementation
Transfer learning is a technique that enables NNs to leverage knowledge acquired from one task to enhance performance on a related but distinct task (refer to Tan et al. 2018 for an overview). This methodology capitalises on pre-existing weight values derived from one data set, employing them as the foundation for network initialisation for training the network on another data set. The advantages of this technique are twofold: it accelerates the training process, and it mitigates network overfitting.
In our work, we utilise as a starting point the Mask R-CNN weights pre-trained on the Microsoft Common Objects in Context (MS COCO; Lin et al. 2014) data set. The MS COCO is a large-scale object detection, segmentation, key-point detection, and captioning dataset. It consists of 328 000328\,000 RGB $3$-channel images of common objects, categorised into $91$ classes. By utilizing transfer learning, we leverage the “universal” low-level geometric features learned from the images in MS COCO (e.g. edges, curves, and corners), allowing our model to generalize effectively despite the relatively small number of instances in our training set. This allows the subsequent training on our specialized dataset to focus exclusively on learning the high-level morphology unique to gravitational arcs. After initialising the pre-trained NN parameters, we made trainable all its $\approx$ 44 million parameters, including also the backbone ones.
@@ 4.2 Performance on the test set
Figure 5: In panel (a) we show two sets of three PPRR curves for $\mathrm{IoU}_{\mathrm{thr}}=0.5,0.75$, @5050:55:$95$, each obtained by varying the object score threshold $p_{\mathrm{thr}}$. In panel (b) we show the PPRR curve for $\mathrm{IoU}_{\mathrm{thr}}=0.5$, where the dots are colour-coded by the score threshold $p_{\mathrm{thr}}$ used to evaluate them. In particular, we emphasise the PPRR pair associated to our choice of score threshold $p_{\mathrm{thr}}=0.996$. Note that the recall remains relatively high even at moderate precision, reflecting the network’s ability to recover most bright arcs while keeping the number of false positives manageable. Increasing the threshold shifts the operating point toward higher precision at the expense of recall. The extended horizontal segments are an artifact of the test set construction: all $500$ test images are simulations based on the same cluster (Abell 1063). As a consequence, a subset of detections (and their associated confidence scores) remains identical across the test set, producing clustered score values and thus plateau-like features in the PR curve. In panel (c) we show two precision and the recall curves as a function of the score threshold for $\mathrm{IoU}_{\mathrm{thr}}=0.5$.
@@ Acknowledgements.
We thank the anonymous referee for the helpful and insightful comments that improved the paper. We acknowledge financial support through grant 2020SKSTHZ. LB is indebted to the communities behind the multiple free, libre, and open-source software packages on which we all depend. MM was supported by INAF Grants The Big-Data era of cluster lensing” and ”Probing Dark Matter and Galaxy Formation in Galaxy Clusters through Strong Gravitational Lensing”, and ASI Grant n. 2024-10-HH.0 ”Attivit scientifiche per la missione Euclid fase E”. The research activities described in this paper have been co-funded by the European Union NextGeneration EU within PRIN 2022 project no. 20229YBSAN Globular clusters in cosmological simulations and in lensed fields: from their birth to the present epoch. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This work used the following software packages: Python (Van Rossum & Drake 2009), PyTorch (Paszke et al. 2019; Ansel et al. 2024), torchvision (TorchVision 2016), NumPy (van der Walt et al. 2011; Harris et al. 2020), SciPy (Virtanen et al. 2020) scikit-learn (Pedregosa et al. 2011) Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), matplotlib (Hunter 2007), lenstool (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009), ds9 (Joye & Mandel 2003), Aladin (Bonnarel et al. 2000), topcat (Taylor 2005), git, bash (GNU 2007).